Parth (parthkohli):

How do you represent the sum of first \(n\) odd numbers through sigma notation?

Parth (parthkohli):

I suppose it is something related to \(2n + 1\) where \(n\) starts at 0?

OpenStudy (anonymous):

\[ \sum_{i=1}^n 2i-1 \]

Parth (parthkohli):

I see.

Parth (parthkohli):

So:\[ \color{Black}{\Rightarrow \sum_{i = 1}^{n} 2i - 1} = n^2\]

Parth (parthkohli):

If it had been \(2i + 1\), then \(i\) would have started at \(0\) and finished at \(n - 1\) right?

OpenStudy (anonymous):

Or you can also take as: \[\large \sum_{k=0}^n 2k+1\]

Parth (parthkohli):

Oh wait. I just got confused.

OpenStudy (anonymous):

Yes, if you start at 0 the sum should only go to n-1.

Parth (parthkohli):

We have a conflict of opinions right there.

Parth (parthkohli):

Let me try it out for \(n = 2\).

Parth (parthkohli):

\[ \sum_{i = 0}^{2 - 1} 2i + 1 = 2(0) + 1 + 2(1) + 1 = 1 + 3\]

Parth (parthkohli):

Yay! I was correct!

Parth (parthkohli):

Thank you @nbouscal :)

OpenStudy (anonymous):

Yes there will come (n-1) sorry..